# Goal-Conditioned Data Augmentation for Offline Reinforcement Learning

### v1

@@ VI Augmented Data Quality Measurement
TABLE IV: Normalized scores of GODA and baseline data augmentation methods. The results are calculated across 5 random seeds. Values in bold represent the best performance (largest score).
Task Dataset IQL TD3+BC
@@
TABLE V: Normalized scores of GODA and baseline data augmentation methods on Adroit tasks evaluated using IQL.
@@
Since our GODA is built upon SynthER, we compare the quality of the datasets augmented by both SynthER and GODA to assess whether the goal conditions incorporated by GODA enhance data quality. We adopt two metrics from SynthER [13]:
Dynamics $\text{Dynamics MSE}=\frac{1}{M}\sum_{i=1}^{M}\left((s_{t+1}^{i},r_{t}^{i})-(\hat{s}_{t+1}^{i},\hat{r}_{t}^{i})\right)^{2},$ (10)
[…]
Average $\text{Average Reward}=\frac{1}{M}\sum_{i=1}^{M}\hat{r}_{t}^{i},$ (12)
Dynamics MSE measures how well the augmentation models capture the dynamics of the environment by learning patterns from the original datasets and generating data that aligns with those dynamics. L2 Distance assesses the models’ exploration capabilities and data diversity by calculating the Euclidean distance between the augmented dataset and the original dataset, reflecting how diverse the generated data is. Average Reward compares the ground-truth reward distributions produced by the simulator given states and actions in datasets augmented by SynthER and GODA.
The left part of Figure 5 presents a scatter plot of 10K points sampled from the augmented datasets. Results show that datasets generated by GODA exhibit much lower Dynamics MSE and a wider range of L2 Distance values, indicating both better alignment with environmental dynamics and greater diversity. The superior validity in performance may stem from the goal conditions (RTG-timestep pairs), which provide critical information for generating samples that better match the environment’s dynamics. Meanwhile, the increased diversity is likely due to the scaled out-of-distribution goal conditions incorporated in the sampling process.
The evaluation results for the three metrics in Table III further demonstrate that GODA outperforms SynthER in terms of all data quality evaluation metrics across nearly all Walker2D tasks.
VII Experimental Results
TABLE VI: Average travel time comparison on real-world TSC tasks. Smaller travel time indicates better traffic efficiency.
@@ VII Experimental Results
VII-A Performance on D4RL
VII-A1 Gym tasks.
Table 1 presents a performance comparison between GODA and other state-of-the-art data augmentation methods trained on the D4RL Gym and Maze2D. We adopt the results of Original and SyntheER from the SynthER paper [13], and those of TATU and DStitch from the DStitch [4] paper. We further conduct experiments for tasks not covered in the literature. As shown in the results, GODA consistently outperforms other methods across most Gym locomotion tasks when evaluated with both IQL and TD3+BC, resulting in higher average scores. Notably, even for tasks using Random datasets, GODA successfully leverages limited high-quality samples to enhance data quality, leading to improved final performance.
VII-A2 Maze2D tasks.
For Maze2D tasks where rewards are sparse, GODA demonstrates significant improvements across all datasets, achieving average gains of 43.4% and 16.0% over the best baseline methods when evaluated with IQL and TD3+BC, respectively. The maximum improvement reaches 57.7% when applying GODA to the Maze2D Large dataset. This highlights GODA’s ability to effectively capture data distributions of various types of datasets and consistently augment high-quality samples.
VII-A3 Adroit tasks.
Table V presents the normalized scores on Adroit Pena and Door tasks, evaluated using the IQL algorithm, as TD3+BC fails to perform on these tasks. As shown, GODA outperforms all baselines on the Pen-Cloned dataset, although it underperforms on Pen-Human. However, for both datasets in the Door task, GODA demonstrates significant improvements over the best baseline methods. These results further demonstrate that GODA is capable of handling more complex tasks.
VII-B Extended Experiments on Traffic Signal Control
@@ VII-C Ablation Study
VII-C1 Ablation on Condition Selection
@@
We test three condition selection mechanisms as described in Section IV-B: return-prior, RTG-prior, and random goal conditions. As shown in the left part of Figure 6, the return-prior method demonstrates superior performance compared to the other two approaches. Moreover, GODA with the return- and RTG-prior conditions outperforms SynthER when tested on two offline RL algorithms. In contrast, the random-prior method shows results comparable to SynthER. This suggests that high-goal conditions identified by the return- and RTG-prior methods enable GODA to generate samples beyond the original data distribution. Randomly selected goal conditions, however, fail to target high-reward regions, producing similar results to SynthER.
VII-C2 Ablation on Goal Scaling Factor
@@
We further examine the effect of different scaling factors on D4RL tasks, testing values ranging from 0.8 to 2.0. As seen in the middle part of Figure 6, the performance improves as the scaling factor increases but slightly degrades when the scaling factor exceeds 1.1. Scaling factors below 1.0 shrink the selected goals, leading to suboptimal samples. Conversely, scaling factors above 1.1 push the selected goals too far beyond the training data distribution, resulting in diminished performance.
VII-C3 Ablation on Adaptive Gated Conditioning
@@
Moreover, we evaluate the impact of the adaptive gated conditioning method. We compare GODA with two variants: one using only adaLN conditioning [33], and another using in-context conditioning, where condition embeddings are directly appended to the input embeddings. From the right part of Figure 6, it is clear that GODA with adaptive gated conditioning achieves the best results and the adaLN and in-context conditioning show similar performance. Additionally, all three methods outperform SynthER which lacks goal conditions. This demonstrates that the inclusion of goal conditions is crucial for guiding the sampling process toward high returns, and our adaptive gated conditioning method enhances the model’s ability to fully utilize these conditions.
VII-C4 Ablation on Top Conditions Selection
@@
VII-C5 Ablation on Mixed Datasets

### v2

@@ 5 Augmented Data Quality Measurement
Table 2: Normalized scores of D4RL Gym and Maze2D tasks for GODA and baseline data augmentation methods. The results are calculated across 5 random seeds. Values in bold represent the best performance (largest score). ‘HC’: HalfCheetah; ‘WK’: Walker2D; ‘HP’: Hopper; ‘MZ’: Maze2D.
Task Data IQL TD3+BC
@@
Table 3: Normalized scores of GODA and baseline data augmentation methods on Adroit tasks evaluated using IQL.
@@
Table 4: Evaluation of GODA on the Cheetah-Run task from the pixel-based V-D4RL benchmark using the BC algorithm, with sample augmentation performed in the latent space.
Eval Algorithm Aug Methods Rand Med-R Med Med-E Average
BC Original 0.0±0.0 25.0±3.6 51.6±1.4 57.5±6.3 33.5±2.8
SynthER 0.1±0.0 27.9±3.4 52.2±1.2 69.9±9.5 37.5±3.5
GODA 1.4±0.8 35.5±6.1 54.9±2.0 67.6±10.4 39.9±4.8
6 Experimental Results
After directly evaluating the data quality of the augmented datasets using various metrics, we further assess GODA by training policies on these datasets and verifying their performance on corresponding tasks. This section presents the overall performance comparison between our GODA and other baseline data augmentation methods.
Table 5: Average travel time comparison on real-world TSC tasks. Smaller travel time indicates better traffic efficiency.
@@ 6 Experimental Results
6.1 Offline Reinforcement Learning Evaluation
6.1.1 Performance on D4RL
Gym tasks. Table 2 presents the normalized score comparison between GODA and other state-of-the-art data augmentation methods trained on the D4RL Gym and Maze2D. We adopt the results of Original and SyntheER from the SynthER paper (Lu et al., 2024), and those of TATU and DStitch from the DStitch (Li et al., 2024) paper. We further conduct experiments for tasks not covered in the literature. As shown in the results, GODA consistently outperforms other methods across most Gym locomotion tasks when evaluated with both IQL and TD3+BC, resulting in higher average scores. Notably, even for tasks using Random datasets, GODA successfully leverages limited high-quality samples to enhance data quality, leading to improved final performance. GODA boosts performance on Med and Med-R datasets; however, a notable gap still remains between the performance of algorithms trained on GODA-augmented Med/Med-R datasets and the performance of algorithms trained on the original Med-E datasets. These limitations suggest that while GODA is effective at maximizing the utility of suboptimal data, it cannot fully synthesize expert-level performance from non-expert data alone. Its strength lies in extracting and amplifying high-quality signals within existing data, not inventing expert behavior from scratch.
Maze2D tasks. For Maze2D tasks where rewards are sparse, GODA demonstrates significant improvements across all datasets, achieving average gains of 43.4% and 16.0% over the best baseline methods when evaluated with IQL and TD3+BC, respectively. The maximum improvement reaches 57.7% when applying GODA to the Maze2D Large dataset. This might be because Maze2D tasks involve navigation from a start to a goal location, similar to the illustrative example in Figure 1, making the goal naturally explicit and spatially defined. GODA can directly condition the learning process on the desired goal position, which aligns perfectly with the task’s structure. Besides, GODA can possibly reuse sub-trajectories and recombine transitions from different paths that end near the same goal, making data augmentation more effective and less likely to violate task constraints. These results highlights GODA’s ability to effectively capture data distributions of various types of datasets and consistently augment high-quality samples.
Adroit tasks. Table 3 presents the normalized scores on Adroit Pena and Door tasks, evaluated using the IQL algorithm, as TD3+BC fails to perform on these tasks. As shown, GODA outperforms all baselines on the Pen-Cloned dataset, although it underperforms on Pen-Human. This might be because Pen task involves high-dimensional, dexterous manipulation, where small perturbations in action can lead to disproportionately large changes in reward (e.g., minor deviations in gripper position can affect pen stability). GODA’s reliance on RTG as a conditioning signal can be limiting in such scenarios. However, for both datasets in the Door task, GODA demonstrates significant improvements over the best baseline methods. These results further demonstrate that GODA is capable of handling more complex tasks.
6.1.2 Performance on V-D4RL
Table 4 shows the results of the BC algorithm evaluated on the pixel-based Cheetah-Run task, with data augmentation performed in the latent space. The results indicate that GODA achieves the best performance on three out of four dataset quality levels and obtains the highest average score. It outperforms the BC algorithm without data augmentation by 19.1%, and the BC with SynthER-augmented samples in the latent space by 6.4%. Notably, GODA delivers substantial improvements on the Random and Medium-Replay datasets, highlighting its suitability for offline scenarios with limited expert demonstrations.
6.1.3 Performance on Traffic Signal Control
@@ 6.2 Ablation Study
6.2.1 Ablation on Condition Selection
@@
We test three condition selection mechanisms as described in Section 3.2: return-prior, RTG-prior, and random goal conditions. As shown in the left part of Figure 5, the return-prior method demonstrates superior performance compared to the other two approaches. Moreover, GODA with the return- and RTG-prior conditions outperforms SynthER when tested on two offline RL algorithms. In contrast, the random-prior method shows results comparable to SynthER. This suggests that high-goal conditions identified by the return- and RTG-prior methods enable GODA to generate samples beyond the original data distribution. Randomly selected goal conditions, however, fail to target high-reward regions, producing similar results to SynthER.
Moreover, return-prior generally outperforms the RTG-prior. This may be because using a return-prior aligns better with the distribution of the original trajectories. However, directly applying the RTG prior can introduce inconsistencies when reordering trajectories, as illustrated in Section 3.2. Specifically, the resulting rewards (calculated as the difference between the RTGs of the current and next timesteps) in the re-ordered trajectories may sometimes lose important high-reward signals, since RTGs are extracted from different original trajectories and may no longer stay in high-reward regions.
6.2.2 Ablation on Goal Scaling Factor
@@
We further examine the effect of different scaling factors on D4RL tasks, testing values ranging from 0.8 to 2.0. As seen in the middle part of Figure 5, the performance improves as the scaling factor increases, but slightly degrades when the scaling factor exceeds 1.1. Scaling factors below 1.0 shrink the selected goals, leading to suboptimal samples. Conversely, scaling factors above 1.1 push the selected goals too far beyond the training data distribution, resulting in diminished performance. Further ablation studies can be seen in Appendix C.4.
Table 6: Ablation study on adaptive gated long skip connection and adaptive gated residual connection. AdaLN removes both the adaptive gated long skip connection and adaptive gated residual connection but only retains the AdaLN structure; w/o AG-LSC removes the adaptive gated long skip connection; w/o AG-RC removes the adaptive gated residual connection; w/o Gate-LSC retains the long skip connection, but without any gating mechanism; and w/o Gate-RC retains the residual connection, but without gating.
Task Eval Alg. AdaLN w/o AG-LSC w/o AG-RC w/o Gate-LSC w/o Gate-RC GODA
WK2D IQL 89.1±5.7 90.2±8.2 94.2±6.1 92.4±8.6 95.0±11.7 96.1±9.4
TD3+BC 91.7±5.9 92.2±4.9 92.5±7.9 91.9±6.5 92.9±9.0 93.0±5.6
Average 90.4±5.8 91.2±6.6 93.4±4.0 92.2±7.6 94.0±10.4 94.6±7.5
6.2.3 Ablation on Adaptive Gated Conditioning
@@
Moreover, we evaluate the impact of the adaptive gated conditioning method. We compare GODA with two variants: one using only adaLN conditioning (Peebles & Xie, 2023), and another using in-context conditioning, where condition embeddings are directly appended to the input embeddings. From the right part of Figure 5, it is clear that GODA with adaptive gated conditioning achieves the best results, and the adaLN and in-context conditioning show similar performance. Additionally, all three methods outperform SynthER, which lacks goal conditions. This demonstrates that the inclusion of goal conditions is crucial for guiding the sampling process toward high returns, and our adaptive gated conditioning method enhances the model’s ability to fully utilize these conditions.
We conduct further ablation experiments on both the condition-adaptive gated long skip connection (Section 3.4.2) and condition-adaptive gated residual connection (Section 3.4.3), and each adaptive gating component as well. The results in Table 6 show that removing either the adaptive gated long skip connection or the adaptive gated residual connection leads to performance degradation, with the adaptive gated long skip connection presenting a more pronounced impact. Replacing the adaptive gated long skip connection with a vanilla long skip connection, i.e., without any adaptive gating, also results in a significant performance drop, though it still performs slightly better than having no long skip connection at all. This highlights the importance of incorporating goal information into the conditioning structure. A similar trend is observed for the adaptive gated residual connection.
6.2.4 Ablation on Top Conditions Selection
@@
While this hyperparameter may seem less critical than $\lambda$, it can still meaningfully influence performance. For example, as shown in Figure 7, the performance difference between the best and worst top-$n$ settings on Walker2D Medium-Replay using IQL reaches 8.8, which is a notable gap. The top-$n$ selection primarily acts as a filter, ensuring that data augmentation focuses on high-return regions. However, increasing $n$ can introduce more diversity, allowing the diffusion model to learn from a broader range of transitions. This represents a classic trade-off between data quality and diversity: smaller $n$ prioritizes high-return samples, while larger $n$ can improve generalization through diversity.
7 Conclusion and Discussion
This paper proposes a novel goal-conditioned data augmentation method for offline RL, namely GODA, which integrates goal guidance into the data augmentation process. We define the easily obtainable return-to-go signal, along with its corresponding timestep in a trajectory, as the goal condition. To achieve high-return augmentation, we introduce several goal selection mechanisms and a scaling method. Additionally, we propose a novel adaptive gated conditioning structure to better incorporate goal conditions into our diffusion model. We demonstrate that data augmented by GODA shows higher quality than SynthER without goal conditions on different evaluation metrics. Extensive experiments on the D4RL benchmark confirm that GODA enhances the performance of classic offline RL methods when trained on GODA-augmented datasets. Furthermore, we evaluate GODA on real-world traffic signal control tasks. The results demonstrate that GODA is highly applicable to TSC problems, even with very small real-world training datasets, making RL-based methods more practical for real-world applications.
We acknowledge that while GODA is effective at maximizing the utility of suboptimal data, it cannot fully synthesize expert-level performance from non-expert data alone. Its strength lies in extracting and amplifying high-quality signals within existing data, not inventing expert behavior from scratch. Besides, GODA can sometimes fail to perform well on some datasets with high reward sensitivity, such as Adroit Pen-Human. Moreover, our current method is not directly applicable to high-dimensional, pixel-based visual RL or real-world robotics datasets, as it requires transforming image data into low-dimensional latent representations before applying diffusion-based data augmentation. Since our approach operates in a low-dimensional feature space, this additional transformation step can complicate the implementation and introduce dependencies on the architecture of the policy network. In future work, we aim to develop more direct methods for augmenting pixel-based datasets by improving the sampling efficiency of the diffusion model. We also plan to explore the application of GODA to more complex, real-world scenarios.
Appendix A Related Work
A.1 Offline Reinforcement Learning
Offline RL involves learning policies from pre-collected offline datasets comprising trajectory rollouts generated by behavior policies. This approach is promising because it avoids the costs and risks associated with direct interactions with the environment. Conventional offline RL methods aim to alleviate the distributional shift problem, i.e., a significant drop in performance due to deviations between the learned policy and the behavior policy used for generating the offline data (Hu et al., 2023). To address this issue, various strategies have been employed, including explicit correction (Xu et al., 2022), such as constraining the policy to a restricted action space (Kumar et al., 2019), and making conservative estimates of the value function (Yu et al., 2021; Kumar et al., 2020), with the aim of aligning the behavior policy with the learned policy.
Some recent studies exploit the strong sequence modeling ability of Transformer models to solve offline RL with trajectory optimization. For instance, Decision Transformer (Chen et al., 2021) and its variants (Wu et al., 2024; Gao et al., 2024) utilize a GPT model to autoregressively predict actions given the recent subtrajectories composed of historical RTGs, states, and actions. These approaches integrate hindsight return information, i.e., RTG, with sequence modeling, eliminating the necessity for dynamic programming.
Diffusion models have also been adopted in offline RL, given their exceptional capability of multi-modal distribution modeling. Diffuser (Janner et al., 2022) employs diffusion models for long-horizon planning, effectively bypassing the compounding errors associated with classical model-based RL. Hierarchical Diffuser (Chen et al., 2024) enhances this approach by introducing a hierarchical structure, specifically a jumpy planning method, to improve planning effectiveness further.
A.2 Data Augmentation in Offline Reinforcement Learning
Rather than passively reusing data and concentrating on enhancing learning algorithms, data augmentation proactively generates more diverse data to improve policy optimization. Some model-based RL methods employ learned world models from offline datasets to simulate the environment and iteratively generate synthetic trajectories, facilitating policy optimization (Zhang et al., 2023). For instance, TATU (Zhang et al., 2023) uses world models to produce synthetic trajectories and truncates those with high accumulated uncertainty. However, model-based RL often suffers from compounding errors in the learned world models. Some basic data augmentation functions (DAFs), i.e., translate, rotate, and reflect, are also applied to augment trajectory segments (Pitis et al., 2020). GuDA (Corrado et al., 2024) further introduces human guidance into these DAFs to improve data quality, while human intervention is costly and lacks scalability. Diffusion models are also directly applied to data augmentation through the sampling process. SynthER (Lu et al., 2024) is the first work that employs diffusion models to learn the distribution of initial offline datasets and unconditionally augment large amounts of new random data. However, it fails to control the sampling process to steer toward high-return directions actively. DiffStitch (Li et al., 2024) attempts to enhance the quality of generated data by actively connecting low-reward trajectories to high-reward ones using a stitching technique.
Figure 7: An illustrative example of how GODA utilizes higher goals to steer the sampling process toward a higher-reward data distribution region.
Appendix B Experiment Details
B.1 Hyperparameters
B.1.1 Denoising Network
Table 7: Hyperparameters for denoiser network.
Hyperparameter Value
embedding dimension 128
MLP width 512
MLP activation SiLU
gate activation Sigmoid
learning rate 0.0003
batch size 256
learning rate schedule cosine annealing
optimizer Adam
gradient update steps 100K
The denoising neural network utilizes the adaptive gated conditioning architecture, as shown in Figure  ref fig:gac. Table 7 details the associated hyperparameters. We use Random Fourier Feature embedding (Rahimi & Recht, 2007) and Sinusoidal positional embedding (Vaswani et al., 2017) to process the noise level and timestep of each RTG, respectively, with an embedding dimension of 128. The width of the linear layers in the MLP block is set to 512, with SiLU (Elfwing et al., 2018) as the activation function. The total number of trainable parameters for the denoiser neural network is approximately 3.3 M. We train our GODA model with 100K steps of gradient updates, with a batch size of 256 and a learning rate of 0.0003.
B.1.2 Elucidated Diffusion Model
Table 8: Hyperparameters for the diffusion model.
Hyperparameter Value
number of diffusion steps 128
$S_{\text{churn}}$ 80
$S_{\text{tmin}}$ 0.05
$S_{\text{tmax}}$ 50
$S_{\text{noise}}$ 1.003
$\sigma_{\text{min}}$ 0.002
$\sigma_{\text{max}}$ 80
We adopt EDM (Karras et al., 2022) as our diffusion model and follow the original settings from SynthER (Lu et al., 2024), with the default hyperparameters shown in Table 8. EDM employs Heun’s $2^{\text{nd}}$ order ODE solver (Ascher & Petzold, 1998) to solve the reverse-time ODE, enabling data sampling through the reverse process. The diffusion timestep is set to 128 for higher-quality results. All training and sampling are conducted on an AMD Ryzen 7 7700X 8-Core Processor and a single NVIDIA GeForce RTX 4080 GPU. Training GODA for 100K steps takes approximately 14 minutes, while generating 5M samples with a sampling batch size of 250K requires about 300 seconds.
B.1.3 D4RL Tasks
In the case of Gym tasks where dense rewards are available, we employ three distinct data configurations from the D4RL datasets: Random, Medium-Replay, and Medium. To elaborate, Random datasets contain transition data generated by a randomly initialized policy. Medium datasets consist of a million data points gathered using a policy that achieves one-third of the performance of an expert approach. Medium-replay datasets contain the stored experience in a replay buffer during the training of a policy until it reaches the score in Medium datasets. Medium-Expert datasets consist of a 50/50 mixture of expert demonstrations and suboptimal trajectories.
For Maze2D, a 2D agent is trained to reach a goal position utilizing minimal feedback, i.e., a single point for success, zero otherwise. Three datasets collected from different maze layouts are adopted, i.e., Umaze, Medium, and Large.
The Pen and Door tasks from the Adroit benchmark (Rajeswaran et al., 2017; Fu et al., 2020) involve manipulating a pen and opening a door using a 24-DoF simulated hand robot. For each task, we use two different datasets: Human and Cloned. The Human dataset consists of trajectories from human demonstrations, while the Cloned dataset is generated by applying an imitation policy trained from a mix of human and expert demonstrations and combining the resulting trajectories with the demonstrations in a 50/50 split.
We augment 5M samples for each D4RL task. For the sampling process, we use the return-prior goal condition selection method and set the goal scaling factor to 1.1 for all tasks. We use the implementation of IQL and TD3+BC from the Clean Offline Reinforcement Learning (CORL) codebase (Tarasov et al., 2024) for D4RL tasks.
B.1.4 TSC Tasks
As shown in Figure 8, a signalized intersection in TSC problems is composed of approaches with several lanes in each approach. The controller manages the phase as shown in the top right part of Figure 8, which determines the activated traffic signals for different directions, to control the orderly movement of vehicles.
((a)) Intersection
((b)) Phases
Figure 8: A standard signalized intersection with four 3-lane approaches. The right part shows eight different phase settings.
We formulate the TSC problem as an MDP and define the state, action, and the reward function as follows:
State. For behavior policies (AMP and ACL), the state representation includes the current phase, traffic movement efficiency pressure, and the number of effective running vehicles (Zhang et al., 2022). For evaluation algorithms, BCQ and CQL use the same state representation as AMP, while DataLight adopts the number of vehicles, along with the total velocity saturation and unsaturation degrees (Zhang & Deng, 2023).
Reward. AMP, BCQ, CQL and DataLight use pressure (Zhang et al., 2022) as the reward while ACL uses queue length. It is worth noting that we use the opposite of these metrics as the final reward function.
We generate a total of 24K samples for each dataset using the behavior policies for each task. Additionally, we augment 20K samples for each task using our GODA model. The task horizon for each TSC scenario is set to 3600 seconds, with a control step of 15 seconds. For the sampling process, we employ the return-prior goal condition selection method and set the goal scaling factor to 1.2 in TSC tasks. For evaluation methods, we employ the implementation of BCQ, CQL, and DataLight from https://github.com/LiangZhang1996/DataLight.
Appendix C Further Experimental Results
In this section, we show some more experimental results for our GODA model.
Table 9: Performance comparison on Gym tasks using DMG as the evaluation algorithm. Values in bold represent the best performance (largest normalized score).
Task Data DMG
Original SynthER GODA
HalfCheetah-v2 Random 27.3±2.5 27.8±1.9 31.5±4.3
Medium-Replay 51.4±1.6 47.1±2.4 50.9±1.7
Medium 55.2±1.8 57.0±0.6 60.5±1.1
Medium-Expert 41.3±2.9 44.2±11.3 46.2±10.6
Walker2D-v2 Random 5.3±1.2 3.6±1.1 13.2±4.9
Medium-Replay 90.9±4.9 91.2±5.1 94.0±4.6
Medium 90.5±5.0 95.7±2.5 88.2±4.7
Medium-Expert 114.0±0.8 115.3±0.5 112.5±0.8
Hopper-v2 Random 9.9±4.4 10.8±3.5 11.2±2.9
Medium-Replay 101.5±1.5 95.4±3.3 99.7±2.8
Medium 100.4±1.0 102.2±0.9 102.2±3.6
Medium-Expert 37.6±33.3 39.1±26.4 36.5±30.1
Average 60.4±5.1 60.8±5.0 62.2±6.0
C.1 Further Evaluation on Gym tasks
To evaluate the generalizability of our proposed method, we further conduct experiments by evaluating GODA with a state-of-the-art offline RL method, i.e., Doubly Mild Generalization (DMG) (Mao et al., 2024), on Gym tasks. DMG introduces a novel generalization strategy by selecting actions from a constrained neighborhood of the training data, thereby maximizing Q-values while avoiding excessive extrapolation. DMG also mitigates generalization propagation and aims to balance generalization and conservatism in offline RL.
As can be seen in Table 9, GODA-enhanced DMG outperforms both the original DMG and the SynthER-enhanced DMG in 7 out of 12 evaluation scenarios, and achieves the highest average normalized score overall. It’s worth noting that we were unable to reproduce the results for HalfCheetah-Medium-Expert and Hopper-Medium-Expert for DMG. SynthER and our GODA also fail to improve the results of DMG on these two datasets to normal levels. We suspect this may be due to hyperparameter sensitivity in DMG rather than issues with the datasets themselves. Overall, these results support the claim that GODA can effectively generalize to more advanced offline RL algorithms and remains a promising data augmentation framework for enhancing state-of-the-art methods.
Algorithm 2 Online Reinforcement Learning with GODA
1:Input: Real-to-synthetic data ratio $\xi$, diffusion model update frequency $\kappa$
2:Initialize policy $\pi_{\phi}$, generative model $G_{\theta}$, real replay buffer $\mathcal{D}^{\text{real}}\leftarrow\emptyset$, synthetic buffer $\mathcal{D}^{*}\leftarrow\emptyset$
3:for $t=1$ to $T$ do
4:  Deploy policy $\pi_{\phi}$ to collect real transitions and store them in $\mathcal{D}^{\text{real}}$
5:  if $t\bmod\kappa=0$ then
6:   Update diffusion model $G_{\theta}$ using samples from $\mathcal{D}^{\text{real}}$ (see Line 2-5 in Algorithm 1)
7:   Generate synthetic samples using $G_{\theta}$ and add them to $\mathcal{D}^{*}$ (see Line 6-11 in Algorithm 1)
8:  end if
9:  Update policy $\pi_{\phi}$ using samples from both $\mathcal{D}^{\text{real}}$ and $\mathcal{D}^{*}$ with a real-to-synthetic ratio of $\xi$
10:end for
Figure 9: Online evaluation with SAC on Gym tasks. GODA significantly outperforms the baselines in the early stages of training, and continues to match or outperform baselines as training progresses.
C.2 Online Evaluation
While our primary focus is on data augmentation for offline RL, our framework can be naturally extended to the online RL setting, similar to the approach taken in SynthER (Lu et al., 2024).
Algorithm 2 shows the pseudocode. In this online extension, we periodically train the diffusion model using transitions collected from ongoing environmental interactions. The learned goal-conditioned diffusion model is then used to augment the dataset with synthetic transitions. The online policy is updated using a hybrid buffer containing both real and synthetic transitions in a 1:1 ratio.
To evaluate this setting, we implemented our method using the Soft Actor-Critic (SAC) (Haarnoja et al., 2018) algorithm and conducted experiments on three Gym tasks: HalfCheetah, Hopper, and Walker2D. We train SAC with an Update-To-Data (UTD) ratio of 1 while training SAC (SynthER) and SAC (GODA) with a UTD of 20. For every 10,000 new real data points gathered, we enhance the dataset by augmenting it with 1 million additional synthetic samples.
Because the distribution and quality of the real buffer improve over time during online training, we adapted our hyperparameters accordingly. Specifically, we applied a decaying goal scaling factor, from 1.8 (effective in low-quality, random offline settings) to 1.1 (effective in expert-level offline settings), and an increasing Top-n condition selection, from 50 to 120, to reflect growing confidence in higher-quality goals as training progresses.
We compare SAC (GODA) with both vanilla SAC and SAC (SynthER), and present online learning curves on Figure 9 showing evaluation returns over 100K interaction steps. As shown, GODA significantly outperforms the baselines in the early stages of training, where real transitions tend to be suboptimal. In the later stages, as the quality and quantity of real data improve, GODA continues to match or slightly outperform SAC (SynthER), with the performance gap gradually narrowing. These results demonstrate that GODA is not only effective in offline settings, but also a promising approach for enhancing online RL through adaptive, goal-conditioned data augmentation.
Table 10: GODA’s sensitivity to dataset complexity. We assess the complexity of offline datasets with several Lipschitz constant metrics and use their 95th percentile values to quantify reward sensitivity. The Pen-Human dataset exhibits the highest sensitivity score, making it particularly challenging for GODA to accurately capture the underlying data distribution and effectively extrapolate to higher-reward regions. Bold values indicate the highest sensitivity.
Dataset LocalLip95 GlobalLip95 GlobalMax RewardStd TrajLip95 GODASensitivity
Door-Cloned 3.361 0.016 5.9 4.0313 7.447 10.16
Door-Human 3.016 1.798 4.8 5.6160 4.757 13.45
Pen-Cloned 29.728 15.896 796.5 28.3944 325.811 106.03
Pen-Human 20.626 9.450 1474.1 29.0234 407.446 109.88
C.3 Failure Case Analysis of GODA
We observe that GODA fails to perform well on some scenarios, such as the Adroit Pen-Human dataset, as shown in Table 3. Therefore, we develop a comprehensive reward sensitivity measurement to examine which offline RL dataset characteristics make GODA effective or problematic. Our method quantifies reward landscape sensitivity using multiple Lipschitz constant metrics across four D4RL Adroit datasets. Local Lipschitz constants are computed using k-nearest neighbors and weighted least squares to estimate how rapidly rewards change with small state-action perturbations at each data point. Global Lipschitz analysis calculate pairwise ratios of reward differences to distances across all neighbor pairs to capture worst-case sensitivity spikes throughout the dataset. Trajectory-level analysis examine within-trajectory sensitivity by computing maximum Lipschitz ratios between trajectory steps. These metrics were combined into a composite GODA Sensitivity weighted by local sensitivity (40%), global sensitivity (30%), reward variability (20%), and trajectory sensitivity (10%), with data normalization and efficient subsampling for computational tractability.
[…]
As can be seen in Table 10, the analysis revealed that task complexity plays an important role on GODA’s success. Door manipulation tasks achieved excellent GODA Sensitivity with low local Lipschitz constants, stable global behavior, and manageable trajectory sensitivity, making them highly preferable for GODA. Conversely, pen manipulation tasks exhibited catastrophic sensitivity with GODA Sensitivity exceeding 100, extreme local Lipschitz values, global sensitivity spikes reaching 800-1500, and volatile trajectory patterns, making them fundamentally challenging for GODA. The critical insight is that the difference between Human and Cloned datasets are minimal within the same task type, but task-intrinsic reward sensitivity creates a significant gap. GODA’s failure on Pen-Human may also stem from the fact that the original dataset contains only 25 human demonstration trajectories, which lack sufficient diversity for GODA to effectively learn the underlying data distribution and extrapolate to out-of-distribution high-reward regions.
C.4 Ablation on Scaling Factor on Different Datasets
To evaluate the practical effectiveness of goal scaling, we conducted additional ablation studies on the Walker2D-Random, Medium, and Medium-Expert datasets.
Figure 10 shows that goal scaling does have a tangible impact: On the Random dataset, performance improved significantly with scaling factors of 1.8 for IQL and 1.4 for TD3+BC. For instance, the score gap between $\lambda=$1.8 and $\lambda=$1.0 in IQL reached 14.6. On Medium and Medium-Replay, a scaling factor of 1.1 consistently yielded the best results for IQL, with gains up to 4.2 compared to $\lambda=$1.0. In contrast, for the Medium-Expert dataset, the difference between $\lambda=$1.0 and $\lambda=$1.1 was marginal. This is likely because datasets with more expert demonstrations have less room for improvement, and higher scaling can push goals outside the feasible return range.
These results suggest that goal scaling is particularly beneficial in low-quality datasets, where it enables exploration of high-return regions beyond what the original data distribution offers. While $\lambda=$1.1 performs robustly across all quality levels and is used as our default, further gains can be achieved through task-specific hyperparameter tuning.
Figure 10: Ablation on scaling factor on different datasets. Goal scaling is particularly beneficial in lower-quality datasets while still necessary for higher-quality datasets.
Figure 11: Ablation on mixed datasets for IQL evaluation.
C.5 Ablation on Mixed Datasets